# JWST captures a sudden stellar outburst and inner disk wall destruction

### v1

@@ 3.1.2 Input parameters
Following previous studies mentioned above, and assuming the outer disk has not changed during previous decades, we take the value of inclination, PA, $\alpha$, $\beta$ and outer radius of the outer disk from Pohl et al. (2017) (see table 2), which is also consistent with recent analysis of ALMA high-resolution continuum data (Hendler et al., 2018). The inner edge of the inner disk is set as 0.04 au ($\sim$5$\times$×R), which is consistent with the optical variable period and estimated to be close to the co-rotation radius $R_{\rm co}$ (e.g., Alcala et al., 1993; Gullbring et al., 1998). The dust temperature at the inner edge is $\sim$ 1800 K from our modelling , which is close to the sublimation temperature of graphite ($\sim$2000 K) but higher than that of silicate ($\sim$1500 K) (Baskin & Laor, 2018). This along with the large gap is in line with the absence of 10 $\mu$m Silicate feature in both Spitzer and JWST spectra. Thus, we set the dust species of the inner disk as purely graphite, while we adopt a dust composition consisting of 70% astronomical silicates and 30% graphite by mass fraction (e.g., Weingartner & Draine, 2001) for the outer disk. We also assume an azimuthally symmetric outer disk well aligned with the inner disk for simplicity.
@@
The inner radius of the outer disk ($r_{\rm in,out}$) has been constrained via NIR interferometry, SED fitting (Olofsson et al., 2011, 2013), near-IR scattered light (Pohl et al., 2017) and millimeter ALMA continuum data (Hendler et al., 2018), and resulting in estimates ranging from 12 au to 30 au. With our model, we found that the inner radius of the outer disk is a crucial parameter and needs to be closer than 24 au so that the outer disk can become hot enough to emit sufficient flux at 15-25 micron even if the inner disk is not blocking any stellar radiation to the outer disk. We allow this parameter to change during our fitting process, and keep it the same between the Spitzer and JWST models. All the parameter samplings are shown in Table 2. Some of the parameters of the inner disk wall, including the $\alpha$, $\beta$, and width ($r_{\rm out,wall}-r_{\rm in,wall}$) do not influence the SED fitting (see appendix A for a detailed discussion). The dust grain size distribution, on the other hand, is highly degenerate with the dust mass because only the small micron-sized grains contribute to the mid-IR SED emission. Thus we take $\alpha$, $\beta$, and the grain size distribution of the inner disk wall the same as Pohl et al. (2017) and keep the $r_{\rm out,wall}=0.06$ au. How the SED changes with each parameter are further discussed in Appendix A.
@@
Note. — Values shown in boldface are first fitted and fixed then between both Spitzer and JWST models. To estimate the upper limit of the mass in the inner disk, the $r_{\rm hidden,out}$ is set at 1 au.
@@ 3.2 Fitted models
By minimizing the chi-square between our model at mid-IR band (5-25 $\mu m$) with the spectra from either Spitzer or JWST, we find a best-fit model for each of them. The results of our best-fit models are summarized in Table 3, while the SEDs compared to the observations are shown in Fig 5. The observed photometric points, spectrum and the input stellar SED are also plotted in Fig 3. The shaded region of the model for the Spitzer spectrum indicates the model with the same parameters but the optically thick part of the inner disk at different azimuthal locations. The dotted pink line indicates the optically thick part is between the star and us, while the dashed line indicate it is away from us. The different azimuthal location of the optically thick wall will also cause a slight difference at mid-IR wavelengths (shaded region in Fig. 5).
@@
Wall Wall Hidden mass Disk without a wall
$(r_{\rm in},r_{\rm out})[au]$ (0.04, 0.06) (0.04, 0.06) (0.06, 1) (0.04, 1)
@@
We note that our asymmetric model is an extreme case with the assumption that all the inner disk mass is azimuthally concentrated in $\delta\phi$ and the whole disk is always in thermal equilibrium. The real mass distribution can be more uniform (i.e., there can be some optically thin or low scale height dust disk in the region where we set the dust mass as 0, contributing to less variable optical extinction than what is assumed in our model). In addition, the cooling timescale of the outer disk ($T_{\rm cool}>0.5$ year, estimated from the equation. 39 of Zhang & Zhu, 2020, with their assumptions and an outer disk inner edge temperature $T\sim 100$K) is much longer than the rotational period of the inner disk ($\sim 3$ days), indicating the outer disk will react less efficiently to the azimuthal location of the inner disk. Thus, the shaded region at both the optical and longer wavelengths ($>10\mu m$) should be interpreted as the maximum variability that can be caused by the inner disk rotation.
@@
As discussed in Section 3 and Olofsson et al. (2013), a large amount of mass can be hidden behind the optically thick inner disk wall. To estimate the upper limit of the mass, we added another component of dust mass behind the inner disk wall with a lower scale height. Considering the inner disk wall height is much lower in JWST time, the maximum mass that can be hidden behind it can be better constrained, we only show the fitted parameters of the additional component in our JWST model. In addition, it is possible that the inner disk wall during the JWST time is sublimated so much that it cannot act as a “wall” that casts enough shadow on the previously hidden mass (see Fig 6 for a sketch). In this case we apply an inner disk with a $r_{\rm in}=0.04$au, $r_{\rm out}=1$au, and other parameters sampled as Table. 2 to fit the JWST spectrum as well. We find that both the inner wall $+$ hidden mass model and the inner disk without a wall model can give a good fit to the JWST spectrum, and the upperlimit of the mass in the inner disk is $\sim 3.6\times 10^{-11}M_{\odot}$, consistent with what is reported in Olofsson et al. (2013) (our value is slightly lower than their estimate $M_{\rm hidden}\sim 5\times 10^{-11}M_{\odot}$). The results of these models are also shown in Table 3.
@@
We find that an azimuthally asymmetric inner disk wall can reproduce the Spitzer spectrum and the strong optical short-term variability of T Cha at the same time. The strong mid-IR seesaw behavior seen in JWST observation along with the higher and less variable optical emission between 2021 and 2023 can be caused by a significant decrease of the inner disk wall height. Due to the degeneracy between parameters, there are also some other models with a slightly higher chi-square that can reproduce a good fit to the data, but all of them follow the same trend as discussed above. Thus we propose that one scenario to explain the phenomenon is that in 2021, an outburst or a stellar magnetic field change caused the inner disk wall to significantly decrease, resulting in the seesaw behavior between the Spitzer and JWST spectra, along with the steady high optical emission. The highly variable optical data and increasing WISE photometry past 2023 suggest that the inner disk wall builds up again.
@@ 4.1 Explanation of our model
In the previous sections, we find that the short-term optical variability seen in T Cha can be caused by an asymmetric inner disk wall. And the significant mid-IR seesaw behavior between Spitzer and JWST spectra can be caused by the inner disk wall decreasing in height or disappearing.
@@
Our model with all the dust mass concentrated in a certain azimuthal region is highly simplified, but a somewhat similar asymmetric inner disk wall can be caused by several mechanisms. At the very inner edge of disks, the dynamical interaction between the magnetic field and the disk can lead to the development of a disk inner wall with azimuthally varying height (e.g., Bouvier et al., 1999). In addition, material from the disk accretes onto the star following the stellar magnetic field lines (e.g., Hartmann et al., 2016). If the accretion originates outside the sublimation radius, the accretion flow can be dusty and thus cause NIR excess while occulting the star periodically (e.g., Bouvier et al., 2007; Gaidos et al., 2024). Both the azimuthally varying disk wall and the dusty accretion flows can act as an azimuthally asymmetric inner disk wall and result in the optical variability we have seen for T Cha. In addition, the perturbation from the gravity of a planet located near the inner disk can also influence the disk dust mass distribution (e.g., Benisty et al., 2023).
@@
We propose that the strong seesaw behavior between Spitzer and JWST spectra along with the stronger but less variable optical emission between 2021 and 2023 is caused by the significant decrease of the inner disk wall height. This decrease is likely due to a stellar outburst happening between 2021-2023, in agreement with the optical emission brightening in that period (see Fig 12 and the blue line and diamonds at optical bands $\sim$2 times higher than the 5600 K photospheric emission in Fig. 3). For the brighter optical emission to be explained simply by less extinction from the inner disk wall, T Cha would require a temperature of at least 6000 K, corresponding to a spectral type earlier than G0 (e.g., Pecaut & Mamajek, 2013). This is significantly different from previous spectra of T Cha showing that it is consistently a G8 type star (e.g., Alcala et al., 1993; Schisano et al., 2009). The discrepancy implies this optical brightening comes from an outburst, and can be tested with future optical spectroscopy.
@@
Since the sublimation radius is very close to the inner edge of the inner disk wall, when the stellar luminosity increases, the dust temperature at the inner disk wall increases and even graphite grains sublimate. In this way, the stellar outburst destroyed the optically thick inner disk dust wall or reduced the wall height significantly, and the strong stellar emission during this period prevents a new tall dust wall to be built. For our model, the innermost asymmetric part has been mostly sublimated and will contribute less to the continuum emission, while the remaining dust which was once hidden behind the inner disk wall (e.g., Olofsson et al., 2013) steadily drifts inwards to maintain the shorter wavelength emission (see Fig 6 for a illustration). This is modelled as a symmetric wall reduced in height and mass plus a hidden mass component, or a inner disk without a wall (see section 3). After 2023 when the star is in a more quiescent stage, the remaining dust moves inward and starts to build up the inner disk wall again.
@@
Such stellar outbursts can be caused by several mechanisms, including the instability of the inner disk edge near the co-rotation radius, and reversal of the stellar magnetic field driven by dynamos (D’Angelo & Spruit, 2010; Armitage, 2016; Fischer et al., 2023). The former scenario has been discussed in detail by D’Angelo & Spruit (2010) and could explain our data as follows. At the quiescent stage, the inner edge of the disk is slightly outside the co-rotation radius where the stellar rotation is faster than the disk. The coupling between the stellar magnetic field and the disk material accelerates the disk material, preventing it from moving inward and suppressing the accretion. During this period, the mass piles up in the inner region of the disk and the inner edge slowly moves inwards, eventually crossing the co-rotation radius. After that, the accretion increases significantly and empties the whole inner region during a short period, causing an accretion outburst while pushing the inner edge of the inner disk outwards, and the system is back to quiescent stage again.
@@ 4.2 Implications of our Model and Future observations
In the previous sections, we proposed that both the optical and mid-IR SED change are caused by a stellar outburst sublimating and destroying most of the inner disk wall. Similar brightening of optical emission was reported in 1989 (Alcala et al., 1993), 2006, and 2015 (see the blue shaded area of Fig. 1), suggesting a recurring event every 5-10 years. This episodic accretion outburst agrees well with the model discussed in D’Angelo & Spruit (2010), and provides several distinct implications.
In the outburst model from D’Angelo & Spruit (2010), during the outburst phase, all the dust mass that crosses the co-rotation radius (i.e., the inner disk wall in our Spitzer model) will be cleared and accreted to the star in a short period. This suggests that every outburst event will decrease the mass of the small (micron-sized) grains in the inner disk by at least $\sim 7\times 10^{-12}M_{\odot}$. Previous studies (Olofsson et al., 2013) and our model from SED fitting suggest that the mass of micron-sized grains in the whole inner disk is less than $3.5\times 10^{-11}M_{\odot}$. On the other hand, with the millimeter flux from Hendler et al. (2018) and subtracting the free-free emission (e.g., Pascucci et al., 2014; Rota et al., 2024), the 3 mm emission from the dust ($\sim 0.1$ mJy) gives a total inner disk dust mass of $\sim 10^{-8}M_{\odot}$. This relatively high dust mass suggests the presence of a lot more millimeter-sized grains compared with micron-sized small grains. From these estimations, we will either be able to see the inner disk of T Cha deplete most of its micron-sized dust grains in the following decades (after $\sim$5 outburst events) which will lead to a significant difference in the SED, or the larger grains will be (or have already been for a long time) colliding with each other to produce second-generation dust, replenishing the loss of micron-sized grains (e.g., Chen et al., 2019; Bernabò et al., 2022). For the latter case, T Cha will be a great place for future studies of second-generation dust and the collision between larger grains. Despite all the previous similar events, we note that our JWST observation is taken at an unusual period, where the system stayed at the possible “outburst” state for a much longer time ($\sim 2$ years compared with previous a few months, see Fig 1). This might also imply that the inner disk is running out of micron-sized dust.
Future observations with different telescopes can provide valuable insights to T Cha while help verify our model. High spatial resolution instruments like VLT/MATISSE can spatially resolve the inner disk of T Cha, which would allow us to further analyze the differences between its outburst and quiescent stages. High-resolution UV or optical spectra taken during the next outburst stage could constrain the accretion rate at that time, verifying whether it is an accretion outburst. Future JWST observations in both quiescent and outburst stage would also reveal whether the spectra are similar to previous observations, while exploring how the disk winds change between stages by analyzing the spectral lines (e.g., Bajaj et al., 2024; Sellek et al., 2024).

### v3

@@ 3.1.2 Input parameters
Following previous studies mentioned above, and assuming the outer disk has not changed during previous decades, we take the value of inclination, PA, $\alpha$, $\beta$ and outer radius of the outer disk from Pohl et al. (2017) (see table 2), which is also consistent with recent analysis of ALMA high-resolution continuum data (Hendler et al., 2018). The inner edge of the inner disk is set as 0.04 au ($\sim$7$\times$×R), very close to the radius corresponding to the $\sim 3.3$ days optical variability. Note that this is a reasonable value given that the magnetic truncation radius is expected to between 0.03-0.06 au (changing the inner edge between these values does not influence the main results, see app. A). With a number density of $\sim 10^{10}$ cm-3 in the inner disk wall and equations 32 and 33 in Baskin & Laor (2018), the sublimation temperature of astronomical silicates is $\sim$1500 K and that of graphite is $\sim$1900 K. Based on the stellar properties, the dust temperature at the inner disk edge is $\sim$ 1800 K, meaning that silicates are sublimated and unlikely to be located as close as 0.04 au while graphite grains can be present.
Thus, we set the dust species of the inner disk wall as purely graphite, while we adopt a dust composition consisting of 70% astronomical silicates and 30% graphite by mass fraction (e.g., Weingartner & Draine, 2001) for the outer disk and the inner disk hidden mass behind the wall. Here we note that the silicates can still be present in the inner disk wall but need to be a minor component as no silicate feature is detected in the Spitzer or JWST spectra. We also assume an azimuthally symmetric outer disk well aligned with the inner disk for both Spitzer and JWST models.
@@
The inner radius of the outer disk ($r_{\rm in,out}$) has been constrained via NIR interferometry, SED fitting (Olofsson et al., 2011, 2013), near-IR scattered light (Pohl et al., 2017) and millimeter ALMA continuum data (Hendler et al., 2018), and resulting in estimates ranging from 12 au to 30 au. With our model, we found that the inner radius of the outer disk is a crucial parameter and needs to be closer than 24 au so that the outer disk can become hot enough to emit sufficient flux at 15-25 µm even if the inner disk is not blocking any stellar radiation to the outer disk. We allow this parameter to change during our fitting process, and keep it the same between the Spitzer and JWST models. The sampling of other parameters is shown in Table 2. Some of the parameters of the inner disk wall, including the $\alpha$, $\beta$, and width ($r_{\rm out,wall}-r_{\rm in,wall}$) do not influence the SED fitting (see appendix A for a detailed discussion). The dust grain size distribution, on the other hand, is highly degenerate with the dust mass because only the small micron-sized grains contribute to the mid-IR SED emission. Thus we take $\alpha$, $\beta$, and the grain size distribution of the inner disk wall the same as Pohl et al. (2017) and keep the $r_{\rm out,wall}=0.06$ au. How the SED changes with each parameter is further discussed in Appendix A.
@@
Note. — Values shown in boldface are first fit and fixed then between both Spitzer and JWST models. To estimate the upper limit of the mass in the inner disk, the $r_{\rm hidden,out}$ is set at 1 au.
@@ 3.2 Fitted models
By minimizing the chi-square between our model at mid-IR wavelengths (5-25 $\mu m$) with the spectra from either Spitzer or JWST, we find a best-fit model for each of them. The results of our best-fit models are summarized in Table 3, while the SEDs compared to the observations are shown in Fig 5. The observed photometric points, spectrum and the input stellar SED are also plotted in Fig 3. The shaded region of the model for the Spitzer spectrum indicates the model with the same parameters but the optically thick part of the inner disk at different azimuthal locations. The dotted pink line indicates the optically thick part is between the star and us, while the dashed line indicate it is away from us. The different azimuthal location of the optically thick wall will also cause a slight difference at mid-IR wavelengths (shaded region in Fig. 5).
@@
Wall Wall Hidden mass
$(r_{\rm in},r_{\rm out})[au]$ (0.04, 0.06) (0.04, 0.06) (0.06, 1)
@@
This is an estimation on the upper limit of the micron-sized dust mass that can be hidden behind the inner disk wall within 1 au.
We note that our asymmetric model is an extreme case with the assumption that all the inner disk mass is azimuthally concentrated in $\delta\phi$ and the whole disk is always in thermal equilibrium. The real mass distribution can be more uniform, i.e. there can be some optically thin or low scale height dust disk in the region where we set the dust mass equal to zero contributing to less variable optical extinction than what is assumed in our model. In addition, the cooling timescale of the outer disk ($T_{\rm cool}>0.5$ year, estimated from equation. 39 of Zhang & Zhu, 2020, with their assumptions and an outer disk inner edge temperature $T\sim 100$K) is much longer than the rotational period of the inner disk ($\sim 3$ days), indicating the outer disk will react less efficiently to the azimuthal location of the inner disk. Thus, the shaded region at both the optical and longer wavelengths ($>10\mu m$) should be interpreted as the maximum variability that can be caused by the inner disk rotation.
@@
As discussed in Section 3 and Olofsson et al. (2013), a large amount of mass can be hidden behind the optically thick inner disk wall. To estimate the upper limit of the mass, we added another component of dust mass behind the inner disk wall with a lower scale height. Considering the inner disk wall height is much lower at the JWST time, hence the maximum hidden mass can be better constrained, we only show the fitted parameters of the additional component in our JWST model. With a good fit to the JWST spectrum, we find an upper limit to the hidden mass in the inner disk of $\sim 3.6\times 10^{-11}M_{\odot}$, consistent with what is reported in Olofsson et al. (2013) (our value is slightly lower than their estimate $M_{\rm hidden}\sim 5\times 10^{-11}M_{\odot}$). The best-fit parameters of this model are reported in Table 3. Besides this, we also explored the possibility that the inner disk wall is completely sublimated and the previously hidden mass is the main contribution to the NIR emission. However, in this case there will be a strong 10 $\mu$m silicate feature, which is not observed in our JWST spectrum (see appendix A for more details). Thus, we can rule out the possibility of complete destruction of the inner wall.
@@
We find that an azimuthally asymmetric inner disk wall can reproduce the Spitzer spectrum and the strong optical short-term variability of T Cha at the same time. The strong mid-IR seesaw behavior seen in the JWST spectrum along with the higher and less variable optical emission between 2021 and 2023 can be caused by a significant decrease of the inner disk wall height. Due to the degeneracy between parameters, there are also some other models with a slightly higher chi-square that can reproduce the data, but all of them follow the same trend as discussed above. Thus we propose that one scenario to explain the observations is that in 2021 an outburst or change in stellar magnetic field caused the inner disk wall to significantly decrease resulting in the seesaw behavior between the Spitzer and JWST spectra, along with the steady high optical emission (see Figure 6 for an illustration). Because the dust sublimation temperature increases with dust density (Baskin & Laor, 2018) and the density of small grains follows the gas density, which has a vertical Gaussian profile, dust in the midplane is less prone to sublimate (while dust above and below sublimates more easily). Doubling the stellar luminosity during the 2021 outburst would only shift the midplane sublimation front to $\sim 0.045$ au, which is why we neglect the outward movement of the inner wall in our analysis. The highly variable optical data and increasing WISE photometry past 2023 suggest that the asymmetric inner disk wall present during the Spitzer observations has built up again.
@@ 4.1 Explanation of our model
In the previous sections, we find that the short-term optical variability seen in T Cha can be caused by an asymmetric inner disk wall while the significant mid-IR seesaw behavior between Spitzer and JWST spectra can be caused by the inner disk wall decreasing in height.
@@
Our model with all the dust mass concentrated in a certain azimuthal region is highly simplified, but a somewhat similar asymmetric inner disk wall can be caused by several mechanisms. At the very inner edge of disks, the dynamical interaction between the magnetic field and the disk can lead to the development of a disk inner wall with azimuthally varying height (e.g., Bouvier et al., 1999). In addition, material from the disk accretes onto the star following the stellar magnetic field lines (e.g., Hartmann et al., 2016). If the accretion originates outside the sublimation radius, the accretion flow can be dusty and thus cause NIR excess while occulting the star periodically (e.g., Bouvier et al., 2007; Gaidos et al., 2024). Both the azimuthally varying disk wall and the dusty accretion flows can act as an azimuthally asymmetric inner disk wall and result in the optical variability (both the periodic one and the significant scatter on top) we have seen for T Cha. In addition, the perturbation from the gravity of a planet located near the inner disk can also influence the disk dust mass distribution (e.g., Benisty et al., 2023).
We propose that the strong seesaw behavior between the Spitzer and JWST spectra along with the stronger but less variable optical emission between 2021 and 2023 is caused by the significant decrease of the inner disk wall height or mass. Recently, Espaillat et al. (2024) analyzed the UX Tau A transition disk showing very similar seesaw behavior at mid-IR wavelengths. The shorter wavelength emission in their JWST spectrum reaches almost the photosphere level while longer wavelength emission increased compared with the Spitzer spectra $\sim 15$ years ago. The contemporary optical photometry is also brighter and less variable. Espaillat et al. (2024) argued that the precession of the inner disk leads to misalignment with the outer disk, disrupting the replenishment of the inner disk. This, along with the constant accretion from the inner disk to the star, depleted the inner disk of micron-sized dust grains, leading to less emission from the inner disk and less shading of the star and the outer disk.
@@
T Cha is similar to UX Tau A except that at the JWST time T Cha has significantly more NIR and optical emission compared with the photosphere emission of a G8 type star even without de-extincting the data. Thus, depleting the whole inner disk is not compatible with the data available for T Cha. We argue, instead that the seesaw behavior of the T Cha spectrum is likely due to a stellar outburst happening between 2021-2023. This explanation agrees with the brightening at optical wavelengths and the stronger ionic lines and PAH features in that period (see Fig 12 and  3 for a reference and Bajaj et al., 2024, for more details). For the brighter optical emission to be explained simply by less extinction from the inner disk wall, T Cha would require a temperature of at least 6000 K, corresponding to a spectral type earlier than G0 (e.g., Pecaut & Mamajek, 2013). This is significantly different from previous spectra of T Cha showing that it is consistently a G8 type star (e.g., Alcala et al., 1993; Schisano et al., 2009). The discrepancy implies this optical brightening can come from an outburst, and can be tested with future optical spectroscopy targeting gas lines.
@@
When the stellar luminosity increases (during the outburst phase), the dust temperature at the inner disk wall increases from $\sim$1800 K to $\sim$2160 K and even graphite grains sublimate. In this way, the stellar outburst could have reduced the wall height significantly, and the strong stellar emission during this period prevents a new tall dust wall to be built. In our model, the innermost asymmetric part has been mostly sublimated and will contribute less to the continuum emission, hence we use a symmetric wall with reduced height and mass (see section 3). After the outburst the hidden dust mass moves inward and starts to build up the tall inner disk wall again.
@@
Such stellar outbursts can be caused by several mechanisms, including the instability of the inner disk edge near the co-rotation radius, and reversal of the stellar magnetic field driven by dynamos (D’Angelo & Spruit, 2010; Armitage, 2016; Fischer et al., 2023). The former scenario has been discussed in detail by D’Angelo & Spruit (2010) and could explain our data as follows. At the quiescent stage, the inner edge of the gas disk (which is at the magnetic truncation radius) is slightly outside the co-rotation radius meaning that the star rotates slightly faster than the inner disk. The coupling between the stellar magnetic field and the disk material accelerates the disk material, preventing it from moving inward, suppressing the accretion, and possibly driving outflows (e.g., Ustyugova et al., 2006). During the quiescent stage, the sublimation radius would be at or inside the truncation radius, both gas and dust would pile up in the inner region of the disk, building up a tall dust wall at the truncation radius. The inner edge slowly moves inwards as mass piles up, eventually crosses the co-rotation radius. After that, the accretion increases significantly and empties the whole inner region during a short period, causing an accretion outburst while pushing the inner edge of the dust disk outward. During this outburst phase, both the corotation radius and the magnetic truncation radius would be inside the sublimation radius -- the dust in the disk will thus cross this radius and sublimates, forming only a lower inner wall at that location due to the pressure maximum by dust sublimation. When the outburst ends, the location of the sublimation radius moves closer to the star, and the system is back to a quiescent stage again. If this scenario is correct, we can also set a constraint on the stellar period of T Cha as shorter than or equal to the $\sim 3.3$ days period as that would set the corotation radius.
@@ 4.2 Implications of our Model and Future observations
In the previous sections, we proposed that both the optical and mid-IR SED changes are caused by a stellar outburst sublimating and destroying most of the inner disk wall. Similar brightening of optical emission was reported in 1989 (Alcala et al., 1993), 2006, and 2015 (see the blue shaded area of Fig. 1), suggesting a recurring event every 5-10 years. This episodic accretion outburst agrees well with the model discussed in D’Angelo & Spruit (2010), and provides several distinct implications.
In the model from D’Angelo & Spruit (2010), during the outburst phase, all the dust mass at the innermost edge of the disk will be sublimated and accreted onto the star in a short period. Based on our model, that would imply that a mass of at least $\sim 7\times 10^{-12}M_{\odot}$ in small grains is removed in 1-2 years. Previous studies (e.g., Olofsson et al., 2013) as well as our model suggest that the mass of micron-sized grains in the whole inner disk is less than $3.5\times 10^{-11}M_{\odot}$. On the other hand, with the millimeter flux from Hendler et al. (2018) and subtracting the free-free emission (e.g., Pascucci et al., 2014; Rota et al., 2024), the 3 mm emission from the dust ($\sim 0.1$ mJy) gives a total inner disk mass of $\sim 10^{-8}M_{\odot}$. The relatively high millimeter dust mass suggests a significantly greater abundance of millimeter-sized grains compared with micron-sized grains. This is also in agreement with the high extinction ratio $R_{V}\sim 5.5$ (e.g., Covino et al., 1997; Schisano et al., 2009) from the disk compared with the average value ($R_{V}=3.1$) for the ISM, which indicates probable grain growth and depletion of small grains (Draine, 2009).
The absence of a gap in the gas disk, along with accretion, suggest that micron-sized grains from the outer disk, which are coupled with the gas, will likely replenish the inner disk. Future ALMA observations and modeling might constrain this replenish rate, providing better insights into on how long the inner micron-sized dust will take to be depleted. Additionally, high spatial resolution instruments like VLT/MATISSE can spatially resolve the inner disk of T Cha, which would allow to further analyze the differences between its outburst and quiescent stages. High-resolution UV or optical spectra taken during the next optical bright and relatively constant stage could constrain the accretion rate at that time, verifying whether T Cha is experiencing an accretion outburst. Future JWST observations in both quiescent and outburst stages would also reveal whether the spectra are similar to previous observations, while also exploring any potential effect from disk winds (e.g., Bajaj et al., 2024; Sellek et al., 2024).